# Internal Consistency and Self-Feedback in Large Language Models: A Survey

### v2

@@ VII Task: Others
Apart from reasoning elevation and hallucination alleviation, several lines of work follow the Self-Feedback framework, although their objectives are not to improve the model’s internal consistency. Examples include knowledge distillation, preference learning, and data augmentation. For the sake of completeness in this survey, this section briefly summarizes these other works adhering to the Self-Feedback framework.
@@ VII-A Preference Learning
Currently, the responses of LLMs in specific domains sometimes fail to satisfy users, even violating human values, social ethics, and morals. The goal of Preference Learning is to enable LLMs to better follow human instructions and output responses that meet human expectations. Most of the work around this task can be broadly covered by the Self-Feedback framework. Specifically, the model generates initial responses during the Self-Evaluation process and then performs Self-Update based on the Feedback Signal. Here, the Feedback Signal mainly refers to the reward information given by a reward model $\mathcal{R}$, which is trained through preference feedback. Preference feedback involves comparing and ranking different responses to the same question in terms of helpfulness, harmlessness, and honesty. The Self-Update here primarily refers to broadly updating the model $\mathcal{M}$, including methods like supervised fine-tuning and reinforcement learning (such as PPO [111], DPO [112]).
There are three main ways to obtain preference feedback. The first is through human feedback, as seen in works like OASST [113] and BeaverTails [114], which include human-annotated data from multiple annotators. The second method involves feedback generated by models [115, 116], offering lower annotation costs and faster iterative feedback efficiency compared to human feedback. The third type of feedback is derived from inductive bias. For instance, the SHP dataset mentioned in [117] considers upvotes and downvotes to differentiate between good and bad responses. Another example is ALMoST [118], which proposes some prior rules, such as larger parameter models generally performing better, and more prompt context examples yielding higher quality responses. These rules are used to rank response quality across different models as preference feedback.
@@ VII-B LLM-Based Knowledge Distillation
Small-parameter LLMs have faster inference and training speeds, but their reasoning capabilities are inferior to those of large-parameter LLMs. LLM-based knowledge distillation methods aim to transfer advanced capabilities from proprietary LLMs (such as GPT-4) to small-parameter open-source models (such as LLaMA and Mistral) [27]. These two models can be referred to as the “teacher model” and the “student model” respectively, with the teacher model guiding the student model to enhance its capabilities, fitting the generalized Self-Feedback framework proposed in this paper. During the Self-Evaluation, the student model generates answers, which are then assessed by the teacher model. In the Self-Update, the student model uses the evaluation signal to update itself or its answers.
This signal can be in the form of statistical metrics, such as MiniLLM [121] calculating the reverse Kullback-Leibler (KL) divergence of the probability distributions output by the student and teacher models; or GKD [122] computing metrics like forward KL divergence, reverse KL divergence, and generalized JSD. The signal can also be in the form of natural language feedback, such as Selfee [123] utilizing ChatGPT as the teacher to provide textual feedback on the outputs of the student model; or in PERsD [77], where the teacher executes the code generated by the student model and provides specific suggestions based on errors.
@@ VII-C Data Augmentation
Currently, methods applying LLMs to Data Augmentation and Synthetic Data Generation mainly focus on the prompt engineering layer. In other words, Self-Evaluation only involves responses. Many studies have shown that LLM responses are highly sensitive to prompt variations [126, 127]. Therefore, the main bottleneck in this task is: how to design better prompts and how to deeply explore the relationship between decoding, latent space, and data quality.
@@ VIII Evaluation
Evaluation helps identify the strengths and weaknesses of different models and methods. This section lists evaluation methods and benchmarks for Internal Consistency and Self-Feedback. These evaluations mainly cover two types of abilities: meta ability, such as model’s uncertainty, consistency, and feedback ability; and common ability, which pertains to solving real-world problems, such as reasoning QA tasks and code generation tasks. Meta evaluation helps us identify which LLMs are more promising for solving complex problems, while common evaluation helps us understand which Self-Feedback methods can better solve problems for a given LLM.
@@ VIII-A Meta Evaluation
We summarize five meta evaluation methods. These methods can generally be categorized into metric-based and benchmark-based paradigms. The former constructs mathematical formulas or metrics to directly calculate the performance of a particular aspect; the latter uses QA datasets to empirically measure performance. We summarize benchmark-based meta evaluations in Table VI.
TABLE VI: Meta Evaluation Benchmarks
@@
Metric-Based Uncertainty Evaluation. As mentioned in Section IV-A, uncertainty estimation involves assessing the uncertainty of a model’s specific response. Uncertainty evaluation, on the other hand, measures the overall uncertainty of a model. Key metrics for evaluating model uncertainty include: Expected Calibration Error (ECE), which assesses the expected difference between model confidence and accuracy; Maximal Calibration Error (MCE), which indicates the maximum deviation between model accuracy and confidence; and Brier Score (BS), which is used to assess how closely the model’s predicted probabilities align with the true class probabilities [25].
Benchmark-Based Uncertainty Evaluation. LLM-Uncertainty-Bench [128] extracts five test tasks (including question answering, reading comprehension, commonsense inference, dialogue response selection, and document summarization) from common benchmark datasets and uses conformal prediction techniques to construct benchmarks. UBench [129] also extracts data from other datasets, totaling 3978 multiple-choice questions covering knowledge, language, understanding, and reasoning abilities. UBench evaluates individual data items by having models textually express uncertainty scores.
Benchmark-Based Consistency Evaluation. This line of work centers on assessing whether a model delivers consistent responses to queries that are semantically equivalent but phrased differently. The key focus is on developing a variety of synonymous queries to test the model’s reliability. For instance, the ConsisEval Benchmark [130] creates simpler synonymous queries for each question. PopQA-TP [131] and ParaRel [132] construct synonymous queries through rephrasing. BMLAMA [133] focuses on multilingual consistency, constructing a parallel corpus of queries. BECEL [134] draws inspiration from behavioral consistency, considering higher-order consistency in model responses by creating semantic consistency data, negational consistency data, symmetric consistency data, etc. Notably, most studies have found that models generally exhibit low consistency.
Benchmark-Based Critique Abilitiy Evaluation. Lin et al. [135] collect a large number of QA pairs from 15 datasets across mathematical, commonsense, symbolic, coding, and algorithmic fields, creating CriticBench through model generation and human annotation. It can be used to evaluate the ability of LLMs to generate critiques, an important aspect of the Self-Feedback framework.
Benchmark-Based Self-Knowledge Evaluation. Self-Knowledge refers to the LLM’s understanding and recognition of its own abilities, limitations, and the content it creates. Yin et al. [4] and Cheng et al. [35] construct sets of unanswerable questions to explore the question “Do large language models know what they do not know?” Tan et al. [136] investigate “Does the model truly understand the questions and solutions it creates?” These studies generally yield negative empirical results, indicating that models have weak Self-Knowledge.
@@ VIII-B Common Evaluation
In most works utilizing the Self-Feedback framework, they do not conduct meta evaluations in experiments but rather assess the model’s ability to solve real-world problems, such as solving math problems, code generation, and summarizing articles. Here, we classify and summarize the popular benchmarks based on the focus of the evaluation tasks [^15] We may see the same benchmark categorized differently in different sources, highlighting the difficulty in achieving fully orthogonal classifications. For example, a dataset evaluating linguistic understanding ability may also involve reasoning tasks. Here, we classify each dataset based on its primary focus., as shown in Table VII. There are two types of reasoning: Knowledge reasoning focuses on the LLM’s ability to solve problems using its parameterized knowledge, while Logic reasoning focuses on complex logical reasoning based on prompts. Linguistic understanding focuses on the LLM’s analysis of language meaning. Code generating and Math Solving respectively test the LLM’s ability to write code and solve math problems.
TABLE VII: Common Evaluation Benchmarks
@@
Currently, these Benchmark evaluation formats mainly include multiple-choice questions and text generation, each with its own advantages and disadvantages. For multiple-choice questions, evaluators can accurately calculate metrics such as accuracy by extracting answer text from the response or comparing the probabilities of different option tokens during LLM decoding. However, many researchers pointed out the drawbacks of this method, namely that LLMs tend to choose the first option, leading to poor self-consistency and not reflecting the LLM’s true ability [143, 144]. Text generation tasks, such as math problems and code generation, mitigate the shortcomings of multiple-choice questions by allowing the LLM to perform freely, better reflecting the model’s true ability. However, this evaluation method struggles to precisely quantify the gap between generated text and reference answers in dimensions such as syntax and semantics. It is evident that current Benchmarks for common evaluation are still far from accurately measuring the true capabilities of LLMs and have many areas for improvement.
@@ IX-A Conflicting Viewpoints
All methods in this survey suggesting that models can generate proper Feedback Signals to optimize themselves. However, with the emergence of a series of works prefixed with “Self-”, questions about their feasibility have also arisen: Can models really self-correct, self refine, etc.?
Jiang et al. [145] propose the SELF-[IN]CORRECT hypothesis. It experimentally verifies in the context of QA tasks that the “accuracy of generating initial answers (generative)” is higher than the “accuracy of judging the correctness of its own generated answers (discriminative)”. This indicates models struggle to assess their own content accurately. They argue that models form a definitive belief about a particular choice during generation, which refutes works like Refining with Responses and Refining the Response Iteratively. However, this work can only refute studies in the domain of QA tasks. Moreover, the model’s discriminative ability is not always greater than its generative ability, so it cannot be concluded that various methods cannot achieve self-correction.
Stechly et al. [146] and Valmeekam et al. [147] used similar strategies to test whether GPT-4 possesses the Self-Feedback capability. Stechly et al. [146] used the Graph Coloring task to test and found that in work similar to Refining the Response Iteratively, GPT-4 almost always fails to verify the correctness of its own solutions but acknowledges methods like Self-Consistency, which suggest that the model can select appropriate solutions from multiple options. In contrast, Valmeekam et al. [147] tested similar strategies in the task planning domain, refuting the feasibility of both Refining the Response Iteratively and Multi-Agent Collaboration. Both Stechly et al. [146] and Valmeekam et al. [147] believe that models have difficulty making correct judgments about their own outputs.
The above criticisms often focus on narrow Feedback Signals. In comparison, the criticisms from Huang et al. [96] are more reasonable. Huang et al. [96] specifically refute the effectiveness of three works (Reflexion [52], Multi-Agent Debate [73], and Self-Refine [8]) through reasonable comparisons and comprehensive experiments. Reflexion relies on external golden truth as a stopping condition for iterative refining during self-correction, which is an unreasonable setup—if there is already a golden truth, there is no need for predictions. For Multi-Agent Debate, the authors’ experiments found that this method is significantly inferior to the Self-Consistency strategy and consumes substantial memory. For Self-Refine, the authors found that the prompts used for initial results and refining were unfair, and they generated a fairer prompt that produced better responses in one go.
@@
Due to the different characteristics of various works, Kamoi et al. [24] provide a more comprehensive analysis by constructing a clear classification method and systematically comparing the strengths and weaknesses of each types of work. They suggest that the ability to self-correct should be discussed according to the specific task. For example, for tasks with decomposable responses [^16] For instance, in the question “Who are some politicians who were born in Boston?”, the feedback generation can be decomposed into easier sub-tasks of verifying each answer. or verifiable tasks [^17] For example, in the Game of 24, a task to find arithmetic operations to obtain 24 using four provided integers, generating a feasible solution might be harder than verification., it is feasible for the model to optimize itself.
@@
Although the survey [24] provides more meaningful viewpoints through classified discussions, it complicates the field, making it difficult to form a systematic framework. Benefiting from the perspective of Internal Consistency and the clear boundary discussions in Section I-E, we conduct a more meaningful discussion on the proposed Self-Feedback framework:
@@ IX-B Does Internal Consistency Mean Correctness?
To answer this question, let’s revisit the relationship between world knowledge, corpus, and language models, as shown in Fig. 9. World knowledge is the consensual (correct) knowledge we humans possess. The training corpus used for models is a true subset of world knowledge, containing the vast majority of correct knowledge and a small portion of uncleanable erroneous knowledge. Additionally, the knowledge embedded in the corpus is deterministic, where each statement in the corpus has a probability of 100%. Language models, by fitting the corpus, acquire higher-order probabilistic representations of this knowledge, and the probabilistic nature makes the learned knowledge vague and non-deterministic, as illustrated by the shaded areas in Fig. 9. Vagueness (or hallucination) is an important characteristic of language models. It enables the generation of novel and creative expressions outside the training corpus distribution. However, from a reliability perspective, vagueness is a disaster. Vagueness means that answers to the same question are uncertain, making the model’s expressions inconsistent.
@@ IX-C Appeals
Currently, there are many chaos in this field: using similarly expressed names, proposing rare or unrealistic tasks, using different benchmarks for the same task, and comparing different baselines. In summary, without clear debate topics, many works contradict each other, resulting in confusing conflicts of views. To avoid these confusions, we propose several appeals.
  • Naming. When proposing new methods, ensure names are appropriate and avoid conflicts (e.g., Self-Improve [46], Self-Improvement [88]). Additionally, when classifying one’s work, consider the accuracy of the direction name. For instance, uncertainty estimation and confidence estimation have different directional indicators; the former leans towards mechanistic explanation, while the latter leans towards application.
  • Task Definition. Some work has begun to notice that reasoning elevation and hallucination alleviation are similar tasks. For example, [148, 149] propose methods that focus on both tasks. We advocate for using Internal Consistency Mining to refer to these two tasks, ensuring terminological standardization.
  • Reasoning and Hallucination. Regarding the use of these two terms, we suggest that when dealing with tasks like QA, we can say that a method may lack reasoning ability; whereas for open-ended generation tasks, we can use exhibit hallucination.
  • Selection of Baselines. This paper summarizes some relatively fixed lines of work, suggesting that when selecting comparison baselines, first determine the sub-direction for your work and choose important work within that direction as comparison baselines. Comparing across different lines of work can lead to unfair comparisons due to significant methodological differences.
  • Experiment Settings. Unrealistic experimental task settings will not advance scientific research. For instance, requiring pre-given golden label to make predictions does not fit real-world needs [96].
  • Prompt Engineering. As mentioned in Section IX-A, many works exhibit a prompt tuning phenomenon, where adjusting the prompt can reverse experimental results. Therefore, we propose that prompt templates must be disclosed in the paper or source code and provide clear usage instructions; verify the robustness of prompts, such as using various templates for experiments; and verify the generality of prompts, such as using multiple different LLMs for experiments.
@@ X-B The Reasoning Paradox
As mentioned in Section II-B, there is a paradox between the reasoning done during single token prediction (latent reasoning [33]) and the reasoning done using multiple tokens in language (explicit reasoning, e.g., CoT).
@@ X-D The Unified Perspective
At present, the focus of work in this field is relatively narrow, lacking a comprehensive understanding of the entire field, and consequently, there are no more general framework works. We believe that using the perspective proposed in this paper, considering problems from the response, decoding, and latent layers in a unified manner, can better facilitate Internal Consistency Mining. There are emerging efforts that begin to integrate multiple layers. For example, Xie et al. [36] start from the response layer and reflects on how different CoT paths guide the consistency of the latent layer; Xie et al. [74] use Self-Evaluation strategies at the response layer to guide better decoding strategies.

### v3

@@ VII Task: Others
Several works follow the Self-Feedback framework, though not always targeting internal consistency. For completeness, we summarize these efforts below.
@@ VII-A Preference Learning
Preference Learning (PL) aims to align LLM outputs with human intent [100, 101, 102]. Most of the work around this task can be broadly covered by the Self-Feedback framework. For PL, the Feedback Signal mainly refers to the reward information given by a reward model $\mathcal{R}$, which is trained through preference feedback. Preference feedback involves comparing and ranking different responses to the same question in terms of helpfulness, harmlessness, and honesty. The Self-Update here primarily refers to broadly updating the model $\mathcal{M}$, including methods like supervised fine-tuning and reinforcement learning (such as PPO [103], DPO [104]).
There are three main ways to obtain preference feedback. 1) Through human feedback, as seen in works like OASST [105] and BeaverTails [106], which include human-annotated data. 2) Feedback generated by models [107, 108], offering lower annotation costs and faster iterative feedback efficiency compared to human feedback. 3) Feedback derived from inductive bias, such as upvotes/downvotes in the SHP dataset [109], or prior rules in ALMoST [110], which rank response quality based on model size or prompt context.
@@ VII-B LLM-Based Knowledge Distillation
LLM-based knowledge distillation methods aim to transfer advanced capabilities from proprietary LLMs (such as GPT-4) to small-parameter open-source models [113]. These two models can be referred to as the “teacher model” and the “student model” respectively, with the teacher model guiding the student model to enhance its capabilities, fitting the generalized Self-Feedback framework proposed in this paper. During the Self-Evaluation, the student model generates answers, which are then assessed by the teacher model. In the Self-Update, the student model uses the evaluation signal to update itself or its answers.
This signal can be in the form of statistical metrics, such as MiniLLM [114] calculating the reverse Kullback-Leibler (KL) divergence of the probability distributions output by the student and teacher models; or GKD [115] computing metrics like forward KL divergence, reverse KL divergence, and generalized JSD. The signal can also be textual feedback, such as Selfee [116] utilizing ChatGPT as the teacher to provide textual feedback on the outputs of the student model; or in PERsD [72], where the teacher executes the code generated by the student model and provides specific suggestions based on errors.
@@ VII-C Data Augmentation
Currently, methods applying LLMs to Data Augmentation and Synthetic Data Generation mainly focus on the prompt engineering layer. In other words, Self-Evaluation only involves responses. Many studies have shown that LLM responses are highly sensitive to prompt variations [119, 120]. Therefore, the main bottleneck in this task is: how to design better prompts and how to deeply explore the relationship between decoding, latent states, and data quality.
@@ VIII Evaluation
This section covers evaluation methods and benchmarks for internal consistency and Self-Feedback, focusing on two abilities: meta (e.g., uncertainty, consistency, feedback) and common (e.g., reasoning QA, code generation) abilities. Meta evaluation identifies which LLMs are the best, while common evaluation reveals which Self-Feedback methods are the best.
@@ VIII-A Meta Evaluation
We summarize five meta evaluation methods, categorized into [Uncaptioned image] metric-based and [Uncaptioned image] benchmark-based approaches. Metric-based methods calculate performance mainly via formulas, while benchmark-based methods empirically measure it using QA datasets (see Table V).
TABLE V: Meta Evaluation Benchmarks
@@
Uncertainty Evaluation [^12] As mentioned in Section IV-A, uncertainty estimation involves assessing the uncertainty of a model’s specific response. Uncertainty evaluation, on the other hand, measures the overall uncertainty of a model.. Key metrics for evaluating model uncertainty include: Expected Calibration Error (ECE), which assesses the expected difference between model confidence and accuracy; Maximal Calibration Error (MCE), which indicates the maximum deviation between model accuracy and confidence; and Brier Score (BS), which is used to assess how closely the model’s predicted probabilities align with the true class probabilities [24].
Uncertainty Evaluation. LLM-Uncertainty-Bench [121] extracts five test tasks (including question answering, reading comprehension, commonsense inference, dialogue response selection, and document summarization) from common benchmark datasets and uses conformal prediction techniques to construct benchmarks. UBench [122] also extracts data from other datasets, totaling 3978 multiple-choice questions covering knowledge, language, understanding, and reasoning abilities. UBench evaluates individual data items by having models textually express uncertainty scores.
Consistency Evaluation. This line of work centers on assessing whether a model delivers consistent responses to queries that are semantically equivalent but phrased differently. The key focus is on developing a variety of synonymous queries to test the model’s reliability. For instance, the ConsisEval Benchmark [123] creates simpler synonymous queries for each question. PopQA-TP [124] and ParaRel [125] construct synonymous queries through rephrasing. BMLAMA [126] focuses on multilingual consistency, constructing a parallel corpus of queries. BECEL [127] draws inspiration from behavioral consistency, considering higher-order consistency in model responses by creating semantic consistency data, negational consistency data, symmetric consistency data, etc. Notably, most studies have found that models generally exhibit low consistency.
Critique Abilitiy Evaluation. Lin et al. [128] collect a large number of QA pairs from 15 datasets across mathematical, commonsense, symbolic, coding, and algorithmic fields, creating CriticBench through model generation and human annotation. It can be used to evaluate the ability of LLMs to generate critiques, an important aspect of the Self-Feedback framework.
Self-Knowledge Evaluation. Self-Knowledge refers to the LLM’s understanding and recognition of its own abilities, limitations, and the content it creates. Yin et al. [4] and Cheng et al. [28] construct sets of unanswerable questions to explore the question “Do large language models know what they do not know?” Tan et al. [129] investigate “Does the model truly understand the questions and solutions it creates?” These studies generally yield negative empirical results, indicating that models have weak Self-Knowledge.
@@ VIII-B Common Evaluation
Self-Feedback methods are often evaluated using benchmarks that focus on real-world tasks like reasoning, code generation, and math problem solving (see Table VI). For more information on LLM evaluation, you can refer to this survey [130].
TABLE VI: Common Evaluation Benchmarks
@@ IX-A Conflicting Viewpoints
With the rise of works prefixed by “Self-”, questions of feasibility arise: Can a model truly optimize itself? Many studies have attempted to answer this question, with most focusing on Refining the Response Iteratively and Multi-Agent Collaboration.
  • Jiang et al. [137] propose the SELF-[IN]CORRECT hypothesis, showing that in QA tasks, models are better at generating answers than judging their own correctness, highlighting a self-assessment limitation.
  • Stechly et al. [138] and Valmeekam et al. [139] found GPT-4 fails to verify its solutions in the Graph Coloring and planning tasks, with verifiers generating many false positives, reducing reliability.
  • Huang et al. [87] refute the effectiveness of Reflexion [48], Multi-Agent Debate [47], and Self-Refine [8]. They argue Reflexion’s reliance on external truth for refining is impractical, Multi-Agent Debate is inferior to Self-Consistency and resource-heavy, and Self-Refine’s prompts were unfair, with better one-shot responses achievable through improved prompting.
@@
  • Kamoi et al. [21] provide a more comprehensive analysis by classifying various methods clearly and systematically comparing the strengths and weaknesses of each methods. They suggest that the ability to self-correct should be discussed according to the specific task. For example, for decomposable tasks [^13] For example, “Who are some politicians who were born in Boston?” or verifiable tasks [^14] For example, in the Game of 24 (Find arithmetic operations to obtain 24 using four given integers), generating a solution is harder than verification., it is feasible for the model to optimize itself.
@@
While these criticisms reveal certain limitations in feedback signals, experimental tasks, and test models, they can be seen as limited perspectives [137, 138, 139, 87]. Although the survey [21] provides more meaningful viewpoints through classified discussions, it complicates the field, making it difficult to form a systematic framework. Benefiting from the perspective of internal consistency and the clear boundary discussions in Section I-E, we conduct a more meaningful discussion on the proposed Self-Feedback framework:
@@ IX-B Does Internal Consistency Mean Correctness?
Let’s revisit the relationship between world knowledge, training corpus, and language models (LMs), as shown in Fig. 10. World knowledge is the consensual (correct) knowledge we humans possess. The training corpus used for models is a true subset of world knowledge, containing the vast majority of correct knowledge and a small portion of uncleanable erroneous knowledge. Additionally, the knowledge embedded in the corpus is deterministic, where each statement in the corpus has a probability of 100%. Language models, by fitting the corpus, acquire higher-order probabilistic representations of this knowledge, but the probabilistic nature makes the learned knowledge vague and non-deterministic, as illustrated by the shaded areas in Fig. 10. Vagueness (or hallucination) is an important characteristic of language models. It enables the generation of novel and creative expressions outside the training corpus distribution. However, from a reliability perspective, vagueness is a disaster. Vagueness means that answers to the same question are uncertain, making the model’s expressions inconsistent.
@@ IX-C Appeals
The field faces significant criticism due to inconsistent naming, unrealistic tasks, varying benchmarks, and contradictory baselines. Thus, we propose the following appeals:
  • Naming. Ensure method names are distinct (e.g., Self-Improve [40] and Self-Improvement [83] are bad names) and accurate (e.g., uncertainty or confidence estimation).
  • Task Definition. Standardize terms by adopting ”Internal Consistency Mining” for reasoning elevation and hallucination alleviation tasks.
  • Reasoning and Hallucination. Use “lack of reasoning” for QA tasks and “exhibiting hallucination” for open-ended generation tasks.
  • Selection of Baselines. Select baselines from the same sub-direction (section) to ensure fair comparisons.
  • Experiment Settings. Avoid unrealistic setups, such as requiring pre-given golden labels [87].
  • Prompt Engineering. Disclose and test prompt templates for robustness and generality across different LLMs.
@@ X-B The Reasoning Paradox
As mentioned in Section II-B, there is a paradox between the reasoning done during single token prediction (latent reasoning [26]) and the reasoning done using multiple tokens in language (explicit reasoning, e.g., CoT) [143].
@@ X-D The Unified Perspective
At present, the focus of work in this field is relatively narrow, lacking a comprehensive understanding of the entire field, and consequently, there are no more general framework works. We believe that using the perspective proposed in this paper, considering problems from the response, decoding, and latent layers in a unified manner, can better facilitate Internal Consistency Mining. There are emerging efforts that begin to integrate multiple layers. For example, Xie et al. [29] start from the response layer and reflect on how different CoT paths guide the consistency of the latent layer; Xie et al. [70] use Self-Evaluation strategies at the response layer to guide better decoding strategies.